OpenStudy (anonymous):

\[g(x) = (1/\sqrt(2\pi) ) e^{-(x-3)^{2}/2}\] find extrema and points of inflection

OpenStudy (anonymous):

Mostly looking to check my answer (got something weird)

OpenStudy (anonymous):

\[g(x) = (1/\sqrt(2\pi) ) e^{-(x-3)^{2}/2}\]

OpenStudy (anonymous):

incase anyone needs the answers: To find extrema find where the first derivative of g(x) is equal to 0. To find the inflections find where the 2nd derivative of g(x) is equal to 0 Max @ (3, \[1/\sqrt(2\pi)\] ) Inflections @ (2\[1/\sqrt(2\pi e)\] ) and (4, \[1/\sqrt(2\pi e )\] )